OpenStudy (anonymous):

4x+5y=-5 3x+y=21 if the system cannot be solved state that it cannot be solved and explain why.. plzz help!!

OpenStudy (turingtest):

My recommendation for this one is to turn the second equation around for y, then sub it into the first equation: 4x+5y=-5 3x+y=21--->y=21-3x ^^^sub this expression for y into the first eqn: 4x+5(21-3x)=-5 once you find x, plug in that value into either eqn to find y ...assuming a unique solution exists

OpenStudy (anonymous):

hmm how can i do that?

OpenStudy (turingtest):

can't solve this for x? 4x+5(21-3x)=-5

OpenStudy (anonymous):

hmm do i multiply 5 with 3x?

OpenStudy (turingtest):

distribute\[a(b+c)=ab+ac\]

OpenStudy (anonymous):

105+15x?

OpenStudy (anonymous):

-15

OpenStudy (turingtest):

yes

OpenStudy (turingtest):

now simplify the whole thing and solve for x

OpenStudy (anonymous):

really i was right? lol

OpenStudy (turingtest):

yeah, and if distribution is new to you then either you have been skipping school or your teacher sucks

OpenStudy (anonymous):

no its not new lol and ive never skip school lol

OpenStudy (turingtest):

I'm just messin' so can you solve for x now?

OpenStudy (anonymous):

yea i know and wait let me see if i can do it lol

OpenStudy (anonymous):

hmm im kinda stuck

OpenStudy (anonymous):

ok i did 4x+105-15x=-5 am i right so far?

OpenStudy (turingtest):

yes

OpenStudy (anonymous):

ok do i put -105 and put it on the -5 and it will be -110?

OpenStudy (turingtest):

yep

OpenStudy (anonymous):

no way! lol

OpenStudy (turingtest):

yes way :)

OpenStudy (anonymous):

so then it will be 11x=-110? am i right i hope so lol

OpenStudy (turingtest):

almost, but watch those negative signs

OpenStudy (anonymous):

o i mean -11 right?

OpenStudy (turingtest):

right, so x=...?

OpenStudy (anonymous):

10?

OpenStudy (turingtest):

right, now we can use any equation we want to find y I suggest remembering our earlier expression we wrote down 4x+5y=-5 3x+y=21--->y=21-3x now that we know x=10, we can find y with the last formula above

OpenStudy (anonymous):

so i plug that into all the x's right?

OpenStudy (turingtest):

you only need to use one of the equations pick whichever you prefer but yes, just sub in x=10 into any of the equations and solve for y

OpenStudy (anonymous):

i did it on the y=21-3(10) and i got y=-9 am i right?

OpenStudy (turingtest):

yes you are :)

OpenStudy (anonymous):

omg so x=10 and y=-9?

OpenStudy (turingtest):

let's make sure they are the right answers our equations were 4x+5y=-5 3x+y=21 if we plug in x=10 and y=-9 both these should be true, or else we messed up

OpenStudy (turingtest):

4(10)+5(-9)=? 3(10)-9=? you should check yourself that these values do in fact satisfy both equations good job :)

OpenStudy (anonymous):

i did it and it came out right=)

OpenStudy (turingtest):

hooray, we win well that covers the three possibilities in systems (none, infinite, or one unique solution), gotta go see ya, and good luck!

OpenStudy (anonymous):

thnx u make me smart lol

OpenStudy (turingtest):

impossible, it must have just been sleeping until now :P

OpenStudy (anonymous):

no u made me smart lol thnx for explaining